■ A. Exercises Match each inequality to the graph of its solution set. 1. (x + Dx 2)<0 2. (x-1)(r +2)20 a. -4-3-2 -1 0 1 2 3 4 -4 -3-2-1 01 2 3 4 -4-3-2-1 0 1 b. C. 4. (x+1 2)>0 5. (x +Dx - 2)s 0 6. (r + 2) <0 7. (x- 1)(r +2)>0 8. (x + 1(x-220 d. e. -4-3 -2-1 0 1 2 3 4 f. g. Please solve and show work/explain thank you!
Answer #1

1) ( x+ 1) ( x- 2) <0

critical points are -1 , 2

testing inequality at each interval

hence, solution is (-1, 2)

1 matches with a

2)  ( x-1) (x+2) >=0

critical points are 1 , -2

testing inequality at each interval

solution is (-infinity , -2 ] U [ 1 , infinity )

2- h

3) ( x- 1) (x+2) <= 0

critical points are 1 , -2

solution is [ -2 , 1 ]

3 - f

4) ( x+1) (x-2) >0

critical points are -1 , 2

solution is ( -infinity , -1 ) U ( 2 , infinity )

4 - c

5)  ( x+1) (x-2) <= 0

solution is [ -1 , 2]

5 - b

6) ( x-1 ) (x+2) < 0

solution is (-2 , 1)

6 - e

7) ( x- 1) (x+2) > 0

7 - g

8) ( x+1) ( x-2) >= 0

8- d